I have to find all the values of b in the following equation that will give one or more real number solutions. Can you please help me with this?
3x^2 + bx - 3 = 0
Thanks!

Answer by stanbon(75887)

I have to find all the values of b in the following equation that will give one or more real number solutions. Can you please help me with this?
3x^2 + bx - 3 = 0
For that to happen the discriminant must be greater than
or equal to zero.
Discriminant = b^2-4*3*(-3)
b^2+36>=0
b^2>=-36
This is true for any Real Number value of "b".
Cheers,
Stan H.